The employment incidence rate e = 15% per week, the unemployment incidence rate u = 0.8%...

  1. The employment incidence rate e = 15% per week, the unemployment incidence rate u = 0.8% per week.
    1. Calculate the equilibrium unemployment rate (assume the constant labour force).
    2. If the government enacted a policy that helped people to find a new job, and this policy raised the employment incidence rate to e = 20% per week, how would the equilibrium unemployment change (give the number (%) and the direction of the change (increase or decrease))?
Answer #1

The formula for computing the unemployment rate is unemployment rate = number of unemployed persons / labor force. Remember that the labor force involves those who are employed plus those who are unemployed.


Equilibrium Unemployment Rate = 0.8/15 * 100= 5.33


The direction of the change (decrease) 0.8/20 * 100 = 4

Please find the below attachment:

Where ASL = Average Supply of Labor, TLF =Total Labor Force, and ADL = Aggregate Demand for Labor.SMTWT Week 44 28 301-064 October 8 9 10 11 12 13 14 15 16 17 18 19 20 21 17 18 19 20 21 22 23 24 25 26 27 28 13 14 15 16 17 1